# On stabilizers in finite permutation groups

### v3

@@ Abstract.
Let $G$ be a permutation group on the finite set $\Omega$. We prove various results about partitions of $\Omega$ whose stabilizers have good properties. In particular, we obtain the following three theorems: In every solvable permutation group, there is a subset whose stabilizer has all orbits of length at most $6$, which is best possible and answers two questions of Babai. Every solvable maximal subgroup of any almost simple group has derived length at most $10$, which is best possible. In every primitive group with solvable stabilizer, there are two points whose pointwise stabilizer has derived length bounded by an absolute constant.
@@ 1. Introduction
In this paper, we are interested in a small variation on these questions. More precisely, we fix only the bare minimum (so one subset, or a pair of points) and we do not aim to find a trivial stabilizer, since this would be impossible in general, but to find a stabilizer that has at least some good properties. We will see that this can be done in a few interesting situations. Regarding subsets, Babai [Bab22] has proved that in every solvable permutation group there exists a set-stabilizer whose orbits have length bounded by a nonexplicit absolute constant. Answering [Bab22, Question 7a], our first main result provides the best possible constant:
@@
Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a solvable permutation group. Then there exists a subset of $\Omega$ whose setwise stabilizer has all orbits of length at most $6$. The constant $6$ is the best possible.
@@
It happens that metabelian stabilizers need not to exist in general (see Example 2.11), so Corollary 1.2 answers [Bab22, Question 7b]. Incidentally, the same example answers negatively a question in [Glu24] on the existence of supersolvable set-stabilizers. We stress that the two constants $6$ and $3$ above turned out to be a little higher than what initially expected, and better results can be obtained for primitive solvable groups. As far as we know, Theorem 1.1 is the first statement concerning stabilizers of imprimitive groups where the property involved is not preserved by a group extension, and part of the difficulty of the proof is in finding the right statement to prove by induction.
With similar ideas, we are able to show that, in every solvable permutation group, stabilizing two disjoint subsets can force the orbits to have length at most $2$. This gives a strong positive answer to [Bab22, Question 7c]. Moreover, we give some elementary proofs concerning nonsolvable groups: we provide an easier proof of a theorem of Cameron [Cam86], and answer [Bab22, Question 6a]. In particular, we prove the following:
@@
Let $G$ be any permutation group on a finite set. Then there exists a setwise stabilizer whose nonalternating composition factors have order bounded by an absolute constant.
@@
We come to the question of fixing two points. Extending the work of Seress [Ser96], Burness [Bur21] showed that in a primitive group with solvable stabilizer, there always exist $5$ points whose pointwise stabilizer is trivial. The hypothesis of having a solvable stabilizer is natural, if we aim to structural results on the stabilizer of two points. For primitive solvable groups, in [LS24] the authors observed that there exist two points whose pointwise stabilizer has bounded derived length. On the other hand, motivated by various applications, Li and Zhang [LZ11] have classified the solvable maximal subgroups of the almost simple groups. Combining their work with some properties of the finite simple groups, we remark the following:
@@
Let $G$ be an almost simple group, and let $H<G$ be a maximal subgroup. If $H$ is solvable, then the derived length of $H$ is at most $10$. The constant $10$ is attained only by the solvable maximal subgroups of the Fischer group $Fi_{{}_{23}}$.
Despite Theorem 1.4, it is easy to exhibit primitive nonsolvable groups of product type with solvable stabilizer of arbitrarily large derived length. Dealing with the product type case, we are able to answer in the affirmative the Question at the end of [LS24]:
@@
The proof of Theorem 1.5 is an application of Corollary 1.2, and it is perhaps surprising that, to answer a question about pointwise stabilizers of primitive nonsolvable groups, one has to deal with set-stabilizers of arbitrary solvable permutation groups. The idea, usually applied in the context of imprimitive linear groups, is using a partition to construct a vector with small centralizer (see [MW04, Sec. 4]). Our argument involves a double coset decomposition and a neat bound gives the constant $13$. It is likely that a more in-depth study of almost simple groups leads to an improvement, but the best possible constant is not even known for solvable groups [LS24].
The paper is organized as follows. In Section 2 we study set-stabilizers, proving Theorem 1.1 and the other related results. In Section 3, we prove Theorems 1.4 and 1.5. The methods are a mixture of group theory and combinatorics.
Acknowledgments: I am grateful to S. Eberhard for many useful conversations. After the public announcement of a first version of this paper, I received precious comments from D. Gluck, H.Y. Huang and S.V. Skresanov. The first asked for a structural result in the spirit of Corollary 2.14 (see also [Glu24]). The second noticed that the (relatively high) constant $10$ in Theorem 1.4 is attained by $Fi_{{}_{23}}$, and that the constant in Theorem 1.5 is at least $5$ for almost simple groups. The third brought the article [Bab22] to my attention, and helped me with the computer calculations in Lemma 2.7. After this, I did a substantial rewrite of the paper and had useful conversations with L. Babai. Finally, I am thankful to L. Pyber for the many ideas he has given me in these years.
@@ 2. Breaking symmetry with set-stabilizers
When $G$ acts on a finite set $\Omega$, then $G$ acts naturally on the power set $\mathcal{P}(\Omega)$. If $\Delta\subseteq\Omega$, we write $C_{G}(\Delta)$ for the pointwise stabilizer, and $Stab_{G}(\Delta)$ for the setwise stabilizer. For each $g\in G$, it is easy to see that the stabilizers of $\Delta$ and $\Delta^{g}$ are conjugate subgroups. As usual, we say that $\Delta$ lies in a regular orbit (i.e., the orbit $\Delta^{G}$) if $Stab_{G}(\Delta)=1$. Of course, as subgroups of $G$, the stabilizers still act on $\Omega$.
@@
For $i\geq 1$, we say that $\Delta\subseteq\Omega$ lies in a $i$-regular orbit if all the orbits of $Stab_{G}(\Delta)$ on $\Omega$ have length at most $i$.
@@
[…]
@@
$Stab_{G}(\Delta)\cong Stab_{G}(\Delta)/C_{G}(\Delta)\leqslant\mathop{\mathrm{Sym}}(b).$
So, if $Stab_{G}(\Delta)$ is solvable, then $dl(Stab_{G}(\Delta))\leq f(b)$, where
@@ 2.1. Primitive solvable groups
The first part follows from [MW04, Lemma 4.1(ii)]. The remaining cases have been handled with the library PrimitiveGroups in [GAP4]. ∎
$\dfrac{1^{1}}{1_{1_{1}}}$ GAP Id Regular orbits on $\mathcal{P}(\Omega)$ $2$-regular orbits on $\mathcal{P}(\Omega)$ $3$-regular orbits on $\mathcal{P}(\Omega)$ Good orbits on $\mathcal{P}(\Omega)$
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $[2,1]$ $1$ $3$ $3$ $3$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[3,1]$ $2$ $2$ $4$ $4$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[3,2]$ 00 $2$ $4$ $4$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[4,1]$ 00 $1$ $3$ $5$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[4,2]$ 00 $1$ $3$ $5$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[5,2]$ 00 $6$ $6$ $8$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[5,3]$ 00 $2$ $2$ $6$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,2]$ $2$ $16$ $16$ $16$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,3]$ $4$ $4$ $10$ $10$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,4]$ 00 $4$ $6$ $8$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[8,1]$ $3$ $5$ $5$ $6$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[8,2]$ 00 00 $3$ $6$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,2]$ 00 $10$ $10$ $24$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,3]$ $4$ $6$ $6$ $14$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,4]$ $4$ $8$ $8$ $12$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,5]$ 00 $4$ $4$ $12$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,6]$ 00 00 $4$ $10$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,7]$ 00 00 $4$ $10$
@@
As before, using [MW04, Lemma 4.1(ii)], and the library PrimitiveGroups in [GAP4]. We observe that $\mathrm{A\Gamma L}_{1}(8)$ and PrimitiveGroup(9,6) do not have any $2$-regular orbit on $\mathcal{P}(\Omega)$, but still have some abelian set-stabilizer. ∎
@@
This is [MW04, Lemma 4.1(iii)]. As before, the proof works by checking the finitely many cases with $|\Omega|\leq 9$. ∎
Lemma 2.4 is sharp, as shown by $\mathop{\mathrm{Sym}}(4)$.
@@ 2.2. Good colorings
The standard argument to deal with imprimitive groups is [Glu83, Theorem 2], which was used by Gluck to show that permutation groups of odd order have a regular orbit on the power set. This does not work with a property such as “have all orbits of length at most $C$”, where $C$ is a constant. We use a different method based on good orbits and partitions that we are going to explain. A key observation, already in the lines of the proof of [Bab22, Theorem 12.1], is that studying stabilizers of partitions can also be useful to obtain results concerning stabilizers of a single subset.
We need some more definitions. An asymmetric coloring of a permutation group $G$ on $\Omega$ is a partition of $\Omega$ with the property that only the identity fixes every part. We refer the reader to [Bab22, Sec. 2] for some history concerning breaking-symmetry results and some conflicting terminology. We refine this notion as follows:
@@
Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, and $k\geq 2$. A $k$-coloring is a coloring of $\Omega$ with at most $k$ colors. For $i\geq 1$, an $i$-asymmetric $k$-coloring is a $k$-coloring such that the stabilizer of the coloring in $G$ has all orbits on $\Omega$ of length at most $i$.
The $1$-asymmetric colorings are precisely the genuine asymmetric colorings. Unless explicitly stated otherwise, the $6$-asymmetric colorings will also be called good colorings. The total number of $k$-colorings of $\Omega$ is $k^{|\Omega|}$, but many of these are equivalent under the action of $G$. Fixed $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, for each $k\geq 2$ we write
[…]
for the numbers of $G$-orbits of asymmetric, $2$-asymmetric, $3$-asymmetric and good $k$-colorings. When necessary, we write $\ell_{k,i}(G)$ to stress the group $G$ we are referring to. It is obvious that, for every $k$ and $i\leq j$,
@@
We observe that describing a $2$-coloring is equivalent to describing a subset. Therefore, $\ell_{2,1}$, $\ell_{2,2}$, $\ell_{2,3}$ and $\ell_{2,6}$ coincide respectively with the numbers of regular, $2$-regular, $3$-regular and good orbits on the power set. Restated in terms of colorings, Lemma 2.4 says that $\ell_{4,1}\geq 1$ for every primitive solvable group (one color is needed for the complement).
@@
The inequality $\ell_{j,i}>\ell_{j-1,i}$ means that there exists a suitable coloring that involves precisely $j$ colors. Now suppose we have such a $j$-coloring. For each subset of cardinality $j$ of the $k$ available colors, we obtain an inequivalent suitable $j$-coloring. ∎
@@
Fix $A$. We have $\ell_{4,1}(A)\geq 1$ by Lemma 2.4. This means that $\ell_{2,1}(A)\geq 1$, or the inequality $\ell_{j,1}(A)>\ell_{j-1,1}(A)$ occurs at some index $j=3,4$. Considering the worst case, we obtain
@@
[…]

### v4

@@ Abstract.
Let $G$ be a permutation group on the finite set $\Omega$. We prove various results about partitions of $\Omega$ whose stabilizers have good properties. In particular, in every solvable permutation group there is a set-stabilizer whose orbits have length at most $6$, which is best possible and answers two questions of Babai. Every solvable maximal subgroup of any almost simple group has derived length at most $10$, which is best possible. In every primitive group with solvable stabilizer, there are two points whose stabilizer has derived length bounded by an absolute constant.
@@ 1. Introduction
In this paper, we are interested in a small variation on these questions. More precisely, we fix only the bare minimum (so one subset, or a pair of points) and we do not aim to find a trivial stabilizer, as that would be impossible in general, but to find a stabilizer that has at least some good properties. Regarding subsets, Babai [Bab22] has showed that in every solvable permutation group there exists a set-stabilizer whose orbits have length bounded by a nonexplicit absolute constant. Answering [Bab22, Quest. 7b], our first main result provides the best possible constant:
@@
Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a solvable permutation group. Then there exists a subset of $\Omega$ whose setwise stabilizer has all orbits of length at most $6$. The constant $6$ is best possible.
@@
It happens that metabelian stabilizers need not exist in general (see Example 2.11), so Corollary 1.2 answers [Bab22, Quest. 7a]. Incidentally, the same example answers negatively a question in [Glu25] on the existence of supersolvable set-stabilizers. We stress that the two constants $6$ and $3$ above turned out to be higher than initially expected, and better results can be obtained for primitive solvable groups. As far as we know, Theorem 1.1 is the first statement concerning stabilizers of imprimitive groups where the property involved is not preserved by a group extension, and part of the difficulty of the proof is in finding the right statement to prove by induction. This strategy, combined with a detailed analysis of primitive groups, turned out to be both powerful and flexible.
With similar ideas, we are able to show that, in every solvable permutation group, stabilizing two disjoint subsets can force the orbits to have length at most $2$. This gives a strong positive answer to [Bab22, Quest. 7c]. Moreover, we give elementary proofs concerning nonsolvable groups: we provide an easier proof of a theorem of Cameron [Cam86], and answer [Bab22, Quest. 6a]. In particular, we prove the following:
@@
Let $G$ be any permutation group on a finite set. Then there exists a set-stabilizer whose nonalternating composition factors have order bounded by an absolute constant.
@@
We come to the question of fixing two points. Extending the work of Seress [Ser96], Burness [Bur21] showed that in a primitive group with solvable stabilizer there always exist $5$ points whose pointwise stabilizer is trivial. The hypothesis of having a solvable stabilizer is natural, if we aim to find structural results on the stabilizer of two points. For primitive solvable groups, in [LS25] the authors observed that there exist two points whose pointwise stabilizer has bounded derived length. On the other hand, motivated by various applications, Li and Zhang [LZ11] have classified the solvable maximal subgroups of the almost simple groups. Combining their work with some properties of the finite simple groups, we remark the following:
@@
Let $G$ be an almost simple group, and let $H<G$ be a maximal subgroup. If $H$ is solvable, then the derived length of $H$ is at most $10$. The constant $10$ is attained only by the solvable maximal subgroups of the Fischer group $\mathop{\mathrm{Fi}}_{{}_{23}}$.
Despite Theorem 1.4, it is easy to exhibit primitive nonsolvable groups of product type with solvable stabilizer of arbitrarily large derived length. Dealing with the product type case, we are able to answer in the affirmative the question at the end of [LS25]:
@@
The proof of Theorem 1.5 is an application of Corollary 1.2 (or of the nonexplicit result of Babai), and it is perhaps surprising that, to answer a question about pointwise stabilizers of primitive nonsolvable groups, one has to deal with set-stabilizers of arbitrary solvable permutation groups. The idea, usually applied in the context of imprimitive linear groups, is using a partition to construct a vector with small centralizer (see the book [MW93] for various examples). Our argument involves a double coset decomposition, and a crude bound gives the constant $13$. It is likely that a more in-depth study of almost simple groups leads to an improvement, but the best possible constant is not even known for solvable groups [LS25, Rem. 2.9].
The paper is simply organized. In Section 2 we study set-stabilizers, proving Theorems 1.1 and 1.3 and other related results. In Section 3, we prove Theorems 1.4 and 1.5. The methods are a mixture of group theory and combinatorics.
Acknowledgments: I am grateful to S. Eberhard for many useful conversations. After the public announcement of a first version of this paper, I received precious comments from D. Gluck, H.Y. Huang and S.V. Skresanov. The first asked for a structural result in the spirit of Corollary 2.14 (see also [Glu25]). The second noticed that the constant $10$ in Theorem 1.4 is attained by $\mathop{\mathrm{Fi}}_{{}_{23}}$, and that the constant in Theorem 1.5 is at least $5$ for almost simple groups. The third brought the article [Bab22] to my attention, and helped me with the computer calculations in Lemma 2.7. After this, I did a substantial rewrite of the paper and had useful conversations with L. Babai. Finally, I am thankful to L. Pyber for the many ideas he has given me in these years.
@@ 2. Breaking symmetry with set-stabilizers
When $G$ acts on a finite set $\Omega$, then $G$ acts naturally on the power set $\mathcal{P}(\Omega)$. If $\Delta\subseteq\Omega$, we write $C_{G}(\Delta)$ for the pointwise stabilizer, and $\mathop{\mathrm{Stab}}_{G}(\Delta)$ for the setwise stabilizer. For each $g\in G$, it is easy to see that the stabilizers of $\Delta$ and $\Delta^{g}$ are conjugate subgroups. As usual, we say that $\Delta$ lies in a regular orbit (i.e., the orbit $\Delta^{G}$) if $\mathop{\mathrm{Stab}}_{G}(\Delta)=1$. Of course, as subgroups of $G$, the stabilizers still act on $\Omega$.
@@
For $i\geq 1$, we say that $\Delta\subseteq\Omega$ lies in a $i$-regular orbit if all the orbits of $\mathop{\mathrm{Stab}}_{G}(\Delta)$ on $\Omega$ have length at most $i$.
@@
 $\mbox{ regular }\Rightarrow\mbox{ 2-regular }\Rightarrow\mbox{ 3-regular }\Rightarrow\mbox{ good }.$
@@
$\mathop{\mathrm{Stab}_{G}}(\Delta)\cong\mathop{\mathrm{Stab}_{G}}(\Delta)/C_{G}(\Delta)\leqslant\mathop{\mathrm{Sym}}(b).$
So, if $\mathop{\mathrm{Stab}}_{G}(\Delta)$ is solvable, then $dl(\mathop{\mathrm{Stab}}_{G}(\Delta))\leq f(b)$, where
@@ 2.1. Primitive solvable groups
The first part follows from [MW04, Lem. 4.1(ii)]. The remaining cases have been handled with the library PrimitiveGroups in [GAP4]. ∎
$\dfrac{1^{1}}{1_{1_{1}}}$ GAP Id Regular orbits on $\mathcal{P}(\Omega)$ 22-regular orbits on $\mathcal{P}(\Omega)$ 33-regular orbits on $\mathcal{P}(\Omega)$ Good orbits on $\mathcal{P}(\Omega)$
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $[2,1]$ 11 33 33 33
$\dfrac{1^{1}}{1_{1_{1}}}$ $[3,1]$ 22 22 44 44
$\dfrac{1^{1}}{1_{1_{1}}}$ $[3,2]$ 0 22 44 44
$\dfrac{1^{1}}{1_{1_{1}}}$ $[4,1]$ 0 11 33 55
$\dfrac{1^{1}}{1_{1_{1}}}$ $[4,2]$ 0 11 33 55
$\dfrac{1^{1}}{1_{1_{1}}}$ $[5,2]$ 0 66 66 88
$\dfrac{1^{1}}{1_{1_{1}}}$ $[5,3]$ 0 22 22 66
$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,2]$ 22 $16$ $16$ $16$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,3]$ 44 44 $10$ $10$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,4]$ 0 44 66 88
$\dfrac{1^{1}}{1_{1_{1}}}$ $[8,1]$ 33 55 55 66
$\dfrac{1^{1}}{1_{1_{1}}}$ $[8,2]$ 0 0 33 66
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,2]$ 0 $10$ $10$ $24$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,3]$ 44 66 66 $14$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,4]$ 44 88 88 $12$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,5]$ 0 44 44 $12$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,6]$ 0 0 44 $10$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,7]$ 0 0 44 $10$
@@
As before, using [MW04, Lem. 4.1(ii)], and the library PrimitiveGroups in [GAP4]. We observe that $\mathrm{A\Gamma L}_{1}(8)$ and PrimitiveGroup(9,6) do not have any 22-regular orbit on $\mathcal{P}(\Omega)$, but still have some abelian set-stabilizer. ∎
@@
This is [MW04, Lem. 4.1(iii)]. ∎
@@ 2.2. Good colorings
The standard argument to deal with imprimitive groups is [Glu83, Th. 2], which was used by Gluck to show that permutation groups of odd order have a regular orbit on the power set. This does not work with a property such as “have all orbits of length at most CC”, where $C$ is a constant. We use a different method based on good orbits and partitions that we are going to explain. A key observation, already in the lines of the proof of [Bab22, Th. 12.1], is that studying stabilizers of partitions can also be useful to obtain results concerning stabilizers of a single subset.
Fix $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, and let $j]=\{1,\ldots,j\}$ (the set of colors). A coloring is a function $\Omega\to[j]$, and a $k$-coloring is a coloring with at most $k$ colors. (This can also be described by the preimages of the colors, which form a partition of $\Omega$.) A coloring is asymmetric if only the identity fixes every color. We refer the reader to [Bab22, Sec. 2] for some history concerning breaking-symmetry results and some conflicting terminology.
@@
For $i\geq 1$, an $i$-asymmetric $k$-coloring is a $k$-coloring such that the stabilizer of the coloring in $G$ has all orbits on $\Omega$ of length at most $i$.
The $1$-asymmetric colorings are precisely the genuine asymmetric colorings. Unless explicitly stated otherwise, the $6$-asymmetric colorings will also be called good colorings. The total number of $k$-colorings of $\Omega$ is $k^{|\Omega|}$, but many of these are equivalent under the action of $G$. Fix $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$; for each $k\geq 2$ we write
$\ell_{k,1},\hskip 14.22636pt\ell_{k,2},\hskip 14.22636pt\ell_{k,3},\hskip 14.22636pt\ell_{k,6},$
for the numbers of $G$-orbits of asymmetric, $2$-asymmetric, 33-asymmetric and good $k$-colorings. When necessary, we write $\ell_{k,i}(G)$ to stress the group $G$ we are referring to. It is obvious that, for every $k$ and $i\leq j$,
@@
We observe that describing a $2$-coloring is equivalent to describing a subset. Therefore, $\ell_{2,1}$, $\ell_{2,2}$, $\ell_{2,3}$ and $\ell_{2,6}$ coincide respectively with the numbers of regular, $2$-regular, 33-regular and good orbits on the power set. Restated in terms of colorings, Lemma 2.4 says that $\ell_{4,1}\geq 1$ for every primitive solvable group (one color is needed for the complement).
@@
The strict inequality $\ell_{j,i}>\ell_{j-1,i}$ means that there exists a suitable coloring that involves precisely $j$ colors. Now suppose we have such a $j$-coloring. For each subset of cardinality $j$ of the $k$ available colors, we obtain an inequivalent suitable $j$-coloring. ∎
@@
Fix AA. We have $\ell_{4,1}(A)\geq 1$ by Lemma 2.4. This means that $\ell_{2,1}(A)\geq 1$, or the inequality $\ell_{j,1}(A)>\ell_{j-1,1}(A)$ occurs at some index $j=3,4$. Considering the worst case, we obtain
@@
$\mathop{\mathrm{Sym}}(2),\hskip 5.69046pt\mathop{\mathrm{Sym}}(3),\hskip 5.69046pt\mathop{\mathrm{Sym}}(4),\hskip 5.69046pt\mathrm{AGL}_{1}(5),\hskip 5.69046pt\mathrm{AGL}_{1}(7),\hskip 5.69046pt\mathrm{A\Gamma L}_{1}(8),\hskip 5.69046pt\mathrm{AGL}_{2}(3).$